What can you say about the returns to scale of the linear production function Q = aK + bL, where a and b are positive constants?
Answer #1

Answer

constant return to scale

-------

if we increase the input n times

output =a(Kn)+b(Ln)

=anK+bnL

=n(aK+bL)

=n(Q)

the output increases by the same time as the input increases, the input increased by n time, and the output increased by n time.